OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 + 4x + 7

OpenStudy (tkhunny):

Complete the Square y = (x^2 + 4x + ______) + 7 - _________

OpenStudy (abb0t):

Start by completing the square.

OpenStudy (anonymous):

i havent learnt this yet

OpenStudy (abb0t):

\[ax^2+bx+c \] to complete the square: \[\left( \frac{ b }{ 2a } \right)^2\]

OpenStudy (anonymous):

oh ok

OpenStudy (abb0t):

and add it to 7

OpenStudy (abb0t):

You should be able to factor to get it in the form: \[y = (x - h)^2 + k\] where (h, k) is the vertex.

OpenStudy (anonymous):

x^2+4x+4+7-4

OpenStudy (abb0t):

Yes. You can finish it from here.

OpenStudy (anonymous):

what form?

OpenStudy (abb0t):

The form I stated above...

OpenStudy (anonymous):

https://www.mathsisfun.com/algebra/completing-square.html D. y =(x + 2)^2 + 3